# Implicit Language Models are RNNs: Balancing Parallelization and Expressivity

### v2

@@ 6 Related Work
The ability of looped models to generalize better to input lengths not seen during training is empirically well established: For example Yang et al. (2024a) show this for looped transformers, while Anil et al. (2022) demonstrate length generalization for DEQs, particularly when they are path independent. Du et al. (2022) show that energy-based models trained to map energy-gradient-descent steps to algorithmic steps, can length generalize in summation, and complex algorithms such as shortest-path. On the theoretical side, The pioneering work of Siegelmann & Sontag (1992) shows that iterated RNNs are Turing complete at infinite numerical precision. More recently, Deletang et al. (2023) studied a number of sequence models and report that grouping tasks by their rung in the Chomsky hierarchy is predictive of models ability to length-generalize. While the works of Merrill et al  (Merrill, 2019; Merrill et al., 2020; Merrill & Sabharwal, 2023; Merrill et al., 2024), which we discuss inSection 2.2, showed that both transformers and SSMs are restricted to TC0; several studies sought to find more precise constraints. Weiss et al. (2021) observe that programs written in a specific language (RASP) can be mapped to transformer models of sufficient capacity. Zhou et al. (2023) then showed that transformers tend to length-generalise if the underlying data-generating process can be expressed in RASP. Sarrof et al. (2024) derived a similar refined constraint for SSMs and showed that they can precisely express star-free regular languages. Grazzi et al. (2024) demonstrate that SSMs can track state in simple problems, such as parity, when their (diagonal) recurrence matrix $\Lambda$ in Equation (1) permits negative eigenvalues. Moreover, they illustrate that a variant of DeltaNet (Yang et al., 2024b) with (possibly) negative eigenvalues can solve the S5 problem when only swaps of two values are considered in the transition. However, no variant of Mamba or DeltaNet was capable of learning S5 and achieving length generalization. To tackle the parallelization-expressiveness trade-off, Beck et al. (2024) propose two new LSTM-inspired layer architectures: the sLSTM and mLSTM layers. While the latter is parallelizable, the former is not and intended to enable the whole model to recognize regular languages. Finally, Soulos et al. (2024) survey strategies for chunking input sequences with transformers, maintaining parallelizability within each chunk and using RNN-like transitions between chunks. They find these architectures recognize regular languages for small chunk sizes with scaling remaining a challenge.
@@ 7 Discussion and Conclusion
Our extensive study of synthetic state tracking problems informs a pretraining schedule for large language models. The implicit Llama and Mamba2 models improve over the baselines in many cases, and prove particularly beneficial on downstream tasks such as HellaSwag (see Table 1). Performance on language modeling is typically primarily determined by parameter count which traditionally caused weight-shared models to underperform (Tay et al., 2023). While implicit models lift the limitations of state-of-the-art language models, self-iteration comes at a cost that only amortizes over the long tail of natural language. However, emerging hardware that accelerates such self-iteration would alleviate this overhead (Brunner et al., 2025). Furthermore, as LLMs make more progress on reducing perplexity, they may eventually face tokens requiring RNN-like transitions.
@@
Finally, given the recent rise of test-time compute (Snell et al., 2024) and latent-space reasoning (Hao et al., 2024), models with adaptive depth per token deserve careful consideration as potential bridgeheads for such techniques as they natively offer adaptive depth and latent-space iteration.
8 Acknowledgment
@@
The authors of the paper would like to thank colleagues from the Analog Optical Computer (AOC) team at Microsoft Research Cambridge for their discussions and feedback during the project. Additionally, we acknowledge support from the Microsoft GCR team for providing the GPUs and prompt assistance in resolving issues faced during the training of large language models.
Appendix A Algebraic Structure of Finite State Machines
@@
This section provides a basic introduction to the word problem and it’s relation to simulating finite state machines (FSMs). We start with some results in the circuit complexity and then relate them to the properties of FSMs.
A.1 Circuit Complexity
@@
A.2 Algebraic Concepts of Finite State Machines
@@
Monoid whose elements can be inverted have a particularly right structure.
@@
Appendix B Proof of Theorem 1
Theorem 1.
Consider an implicit SSM given by Equations (7) and (6). Then the transition function $h_{t-1}^{*}\mapsto h_{t}^{*}$ is non-linear and non-diagonal, i.e. each hidden state $h_{t}^{*}$ is a non-linear function of the previous hidden state $h_{t-1}^{*}$. Consequently, the state-to-state Jacobian is a non-diagonal operator.
Proof.
We will apply the implicit function theorem for the function
$g\left(z,h,x,\theta\right)=z-f_{\theta}\left(z,h,x\right)\,.$ (10)
If $g\left(z_{t}^{*},h_{t-1}^{*},x_{t},\theta_{0}\right)=0$ and $J_{g,z}$ is non-singular, then there exists an open set $U$ with $\left(h_{t-1}^{*},x,\theta_{0}\right)\in U$ and a differentiable function $\varphi$ on $U$ s.t.
The derivative of $\varphi$ at the fixed point is given by
$\left.\frac{\partial\varphi}{\partial h}\right|_{h_{t-1}^{*},x_{t},\theta_{0}}=-\left(I-\left.\frac{\partial f_{\theta}}{\partial z}\right|_{z_{t}^{*},h_{t-1}^{*},\theta_{0}}\right)^{-1}\left.\frac{\partial f_{\theta}}{\partial h}\right|_{z_{t}^{*},h_{t-1}^{*},\theta_{0}}$ (12)
Clearly, $\varphi$ is a non-linear function if $f_{\theta}$ is a non-linear function.
Now, consider Equation 6 at the fixed point
$\left.\frac{\partial h_{t}^{*}}{\partial h_{t-1}^{*}}\right|_{h_{t-1}^{*},x_{t},\theta}$ $=\Lambda\left(z_{t}^{*},x_{t}\right)$
$+\frac{\partial\Lambda}{\partial z_{t}^{*}}\frac{\partial\varphi}{\partial h_{t-1}^{*}}\mathrm{diag}\left(h_{t-1}^{*}\right)$
This equation highlights the non-diagonal corrections to the diagonal Jacobian $\Lambda$ of the explicit state-space Equation (1). ∎
Figure 5: Fixed-point iteration and phantom gradients: A neural network is iterated until convergence in the forward pass. When employing the phantom gradient principle, only a fraction of the forward steps is however considered for the backward pass.
Appendix C Additional Results
C.1 Inference Dynamics and Convergence of Implicit Language Models
Model D-Pile Perplexity Inference Steps Rel. Diff.
130M Mamba2(4+1) 13.76 14 0.035
Mamba2(24+4) 12.86 62 0.036
Llama(4+1) 12.73 13 0.014
Llama(32+4) 11.73 53 0.033
350M Mamba2(4+1) 10.02 10 0.012
Mamba2(24+4) 9.70 49 0.024
Llama(4+1) 9.66 12 0.015
Llama(32+4) 9.43 57 0.037
760M Mamba2(4+1) 8.60 10 0.013
Mamba2(24+4) 8.35 45 0.025
Llama(4+1) 8.27 12 0.014
Llama(32+4) 7.90 77 0.044
1.3B Mamba2(4+1) 7.97 10 0.013
Mamba2(24+4) 7.70 47 0.029
Llama(4+1) 7.66 13 0.015
Llama(32+4) 7.24 69 0.048
C.2 Length Extrapolation Capabilities in Pretrained Models
Model 2048 4096 8192 16384
$\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$
130M Mamba2 13.72 13.44 13.92 14.91
Mamba2 13.05 12.72 12.76 12.94
Mamba2(4+1)-ours 13.76 13.04 13.06 13.25
Mamba2(24+4)-ours 12.86 12.53 12.55 12.78
350M Mamba2 10.55 10.28 11.20 13.89
Mamba2 10.18 9.94 10.07 10.34
Mamba2(4+1)-ours 10.02 9.74 9.73 9.86
Mamba2(24+4)-ours 9.70 9.45 9.44 9.59
760M Mamba2 9.23 9.45 34.99 231.25
Mamba2 8.98 8.94 10.38 12.56
Mamba2(4+1)-ours 8.60 8.38 8.40 8.55
Mamba2(24+4)-ours 8.35 8.16 8.18 8.33
1.3B Mamba2 8.40 8.47 12.69 25.37
Mamba2 8.28 8.24 9.96 15.06
Mamba2(4+1)-ours 7.97 7.81 8.36 9.53
Mamba2(24+4)-ours 7.70 7.59 8.19 9.63
C.3 Effective Duality between Simultaneous Mode and Sequential Mode in SSMs
Ground Truth Simultaneous Sequential
Appendix D Experimental Details
D.1 The Word Problem
Details of the Learning Problem
@@ D.2 CatbAbI
The models, both implicit and explicit, comprise up to three layers of the Mamba2 architecture with an embedding dimension of 256, a state dimension of 16 (expansion factor 2), and a head dimension of 32. We trained the models using batch sizes of 128 and 256, and learning rates of 0.0001, 0.0005, 0.001, and 0.005. The models were trained for 15,000 steps, with the implicit model specifically trained for 5,000 steps in unrolling mode, utilizing 32 steps with normal gradient checkpointing, followed by 10,000 steps of self-iteration fixed-point search. The self-iteration included a stop threshold of 0.03 and a training and testing maximum number of steps 50 and 200 , respectively, and phantom gradient parameters of 6 steps with ($\lambda=0.5$). Data were packed in sequences of length 200 tokens as per (Schlag et al., 2021). Figures 7 and 8 illustrate the validation accuracy of the explicit and implicit Mamba2 models on the CatbAbI dataset, respectively. Additionally, Figure 9 plots the number of iterations required for the implicit model to reach a fixed point on the validation set of the CatbAbI dataset.

### v3

@@ 6 Related Work
The ability of looped models to generalize better to input lengths not seen during training is empirically well established: For example Yang et al. (2024a) show this for looped transformers, while Anil et al. (2022) demonstrate length generalization for DEQs, particularly when they are path independent. Du et al. (2022) show that energy-based models trained to map energy-gradient-descent steps to algorithmic steps, can length generalize in summation, and complex algorithms such as shortest-path. On the theoretical side, The pioneering work of Siegelmann & Sontag (1992) shows that iterated RNNs are Turing complete at infinite numerical precision. More recently, Deletang et al. (2023) studied a number of sequence models and report that grouping tasks by their rung in the Chomsky hierarchy is predictive of models ability to length-generalize. While the works of Merrill et al  (Merrill, 2019; Merrill et al., 2020; Merrill & Sabharwal, 2023; Merrill et al., 2024), which we discuss inSection 2.2, showed that both transformers and SSMs are restricted to TC0; several studies sought to find more precise constraints. Weiss et al. (2021) observe that programs written in a specific language (RASP) can be mapped to transformer models of sufficient capacity. Zhou et al. (2023) then showed that transformers tend to length-generalise if the underlying data-generating process can be expressed in RASP. Sarrof et al. (2024) derived a similar refined constraint for SSMs and showed that they can precisely express star-free regular languages. Grazzi et al. (2025) demonstrate that SSMs can track state in simple problems, such as parity, when their (diagonal) recurrence matrix $\Lambda$ in Equation 1 permits negative eigenvalues. Moreover, they illustrate that a variant of DeltaNet (Yang et al., 2024b) with (possibly) negative eigenvalues can solve the S5 problem when only swaps of two values are considered in the transition. However, no variant of Mamba or DeltaNet was capable of learning S5 and achieving length generalization. To tackle the parallelization-expressiveness trade-off, Beck et al. (2024) propose two new LSTM-inspired layer architectures: the sLSTM and mLSTM layers. While the latter is parallelizable, the former is not and intended to enable the whole model to recognize regular languages. Finally, Soulos et al. (2024) survey strategies for chunking input sequences with transformers, maintaining parallelizability within each chunk and using RNN-like transitions between chunks. They find these architectures recognize regular languages for small chunk sizes with scaling remaining a challenge.
@@ 7 Discussion and Conclusion
Our extensive study of synthetic state tracking problems informs a pretraining schedule for large language models. The implicit Llama and Mamba2 models improve over the baselines in many cases, and prove particularly beneficial on downstream tasks such as HellaSwag (see Table 1). Performance on language modeling is typically primarily determined by parameter count which traditionally caused weight-shared models to underperform (Tay et al., 2023). While implicit models lift the state-tracking limitations of explicit language models, state-of-the-art explicit models perform extraordinarily well on natural language. Self-iteration introduces additional cost that might only amortize on specific problems requiring state-tracking over long sequences. However, emerging hardware that accelerates such self-iteration would alleviate this overhead (Brunner et al., 2025). Furthermore, as LLMs make more progress on reducing perplexity, they may eventually face tokens requiring RNN-like transitions.
@@
Finally, given the recent rise of test-time compute (Snell et al., 2025) and latent-space reasoning (Hao et al., 2024), models with adaptive depth per token deserve careful consideration as potential bridgeheads for such techniques as they natively offer adaptive depth and latent-space iteration.
Impact Statement
This paper presents work aimed at advancing the field of machine learning by developing sequential and language models that can track state and hence prove to be more powerful in reasoning. Regarding language modeling, there are numerous potential societal consequences to consider, such as the energy consumption required for the training of large models and the additional iterations that our implicit models impose (see Table 4) to enhance computational expressivity of language models. When developing training algorithms for our language modeling, computational efficiency was taken into account— see Section D.3. Nevertheless, further research is essential to maintain the power of the proposed models while continuing to reduce computational costs.
Acknowledgment
@@
The authors of the paper would like to thank colleagues from the Analog Optical Computer (AOC) team at Microsoft Research Cambridge for their discussions and feedback during the project. Additionally, we acknowledge support from the Microsoft GCR team for providing the GPUs and prompt assistance in resolving issues faced during the training of large language models. MS was partially supported with funds from Bosch-Forschungsstiftung im Stifterverband.
Appendix A Implicit State-Space Models
A.1 Fixed Point Iteration
Figure 5 visualizes the forward and backward pass of implicit models. A fixed point search is conducted by unrolling the model until the relative difference $\frac{\lvert z_{t}^{\left(s\right)}-z_{t}^{\left(s-1\right)}\rvert}{\lvert z_{t}^{\left(s-1\right)}\rvert}$ falls below a tolerance $\epsilon$, where we choose $\epsilon=0.05$ for the language models and $\epsilon=0.01$ for the state-tracking models.
Figure 5: Fixed-point iteration and phantom gradients: A neural network is iterated until convergence in the forward pass. When employing the phantom gradient principle, only the final $k$ steps of the forward pass are considered for the backward pass as expressed in Equation 5. Therefore, the memory required for the backward pass is proportional to $k$ and in particular not proportional to the number of steps in the forward pass.
The phantom gradient approach backpropagates only through a fixed number of steps, effectively decoupling the number of iterations in the forward pass from the number of iterations in the backward pass.
Figure 6 shows the two loops over the sequence $t$ and over depth $s$ as presented in Section 3.1. The equations involving the iteration variables are printed again below.
$h_{t}^{\left(s\right)}$ $=\Lambda\left(z_{t}^{\left(s-1\right)},x_{t}\right)h_{t-1}^{\left(s\right)}+u\left(z_{t}^{\left(s-1\right)},x_{t}\right)$
$z_{t}^{\left(s\right)}$ $=f_{\theta}\left(z_{t}^{\left(s-1\right)},h_{t-1}^{\left(s\right)},x_{t}\right)\,.$
A.2 Proof of Theorem 1
Theorem 1 (Extended).
An implicit SSM defined by Equations 8 and 9 with generic weights yields a non-linear and non-diagonal state-to-state transition function $\xi:h_{t-1}^{*}\mapsto h_{t}^{*}$. Let ${g\left(z,h,x,\theta\right)=z-f_{\theta}\left(z,h,x\right)}$. If $g\left(z_{t}^{*},h_{t-1}^{*},x_{t},\theta\right)=0$ and ${J_{g,z}|_{h_{t-1}^{*},x_{t},\theta}=\left.\frac{\partial g}{\partial z}\right|_{h_{t-1}^{*},x_{t},\theta}}$ is non-singular, there exists an open set $U$ with $\left(h_{t-1}^{*},x_{t},\theta\right)\in U$ and a continuously differentiable function $\varphi$ on $U$ s.t.
$g\left(\varphi\left(h,x,\theta\right),h,x,\theta\right)=0\quad\forall\left(h,x,\theta\right)\in U\,.$
In this case, the state-to-state Jacobian $J_{\xi}$ is given by
$\left.\frac{\mathrm{d}h_{t}^{*}}{\mathrm{d}h_{t-1}^{*}}\right|_{z_{t}^{*},x_{t},\theta}=\Lambda\left(z_{t}^{*},x_{t}\right)+\frac{\partial\Lambda}{\partial z_{t}^{*}}\frac{\partial\varphi}{\partial h_{t-1}^{*}}\mathrm{diag}\left(h_{t-1}^{*}\right)+\frac{\partial u}{\partial z_{t}^{*}}\frac{\partial\varphi}{\partial h_{t-1}^{*}}\,.$
Proof.
Suppose that fixed points $h_{t-1}^{*},z_{t-1}^{*}$ have been found for $t-1$. Our goal is to back the above statements for $t$. Let $T\left(z\right)=f_{\theta}\left(z,h_{t-1}^{*},x_{t}\right)$. With generic weights $\theta$, e.g. $\theta$ in general linear position, a multi-layer neural network $f_{\theta}$ is non-linear and its input-output function $T\left(z\right)$ has a non-diagonal Jacobian. Finding the fixed point $z_{t}^{*}$ can be expressed in terms of recursive application of the non-linear operator $T$ as
$z_{t}^{*}=\lim_{s\rightarrow\infty}z_{t}^{\left(s\right)}=\lim_{s\rightarrow\infty}T^{s}\left(z_{t}^{\left(0\right)}\right)\,,$ (11)
where we set $z_{t}^{\left(0\right)}=0$. By extension from multi-layer networks to self-iterated multi-layer networks, generic weights yield a non-linear function ${\left(h_{t-1}^{*},x_{t},\theta\right)\longrightarrow z_{t}^{*}}$. Once the fixed point $z_{t}^{*}$ is found, we can compute $h_{t-1}^{*}$ by a single pass through Equation 8.
To proof Equation 10, we apply the implicit function theorem to the function
$g\left(z,h,x,\theta\right)=z-f_{\theta}\left(z,h,x\right)\,.$ (12)
If $g\left(z_{t}^{*},h_{t-1}^{*},x_{t},\theta\right)=0$ and ${J_{g,z}|_{h_{t-1}^{*},x_{t},\theta}=\left.\frac{\partial g}{\partial z}\right|_{h_{t-1}^{*},x_{t},\theta}}$ is non-singular, the implicit function theorem guarantees the existence of an open set $U$ with $\left(h_{t-1}^{*},x,\theta\right)\in U$ and a continuously differentiable function $\varphi$ on $U$ s.t.
The derivative of $\varphi$ at the fixed point is given by
$\left.\frac{\partial\varphi}{\partial h}\right|_{h_{t-1}^{*},x_{t},\theta}=\left(I-\left.\frac{\partial f_{\theta}}{\partial z}\right|_{z_{t}^{*},h_{t-1}^{*},x_{t},\theta}\right)^{-1}\left.\frac{\partial f_{\theta}}{\partial h}\right|_{z_{t}^{*},h_{t-1}^{*},x_{t},\theta}$ (14)
By the above argument, $\varphi$ is a non-linear function if $f_{\theta}$ is a non-linear function, and ${\varphi\left(h_{t-1}^{*},x_{t},\theta\right)=z_{t}^{*}=\lim_{s\rightarrow\infty}T^{s}\left(0\right)}$.
Now, consider Equation 8 at the fixed point
$\left.\frac{\mathrm{d}h_{t}^{*}}{\mathrm{d}h_{t-1}^{*}}\right|_{z_{t}^{*},x_{t},\theta}$ $=\Lambda\left(z_{t}^{*},x_{t}\right)\frac{\partial h_{t-1}^{*}}{\partial h_{t-1}^{*}}+\frac{\partial\Lambda}{\partial h_{t-1}^{*}}h_{t-1}^{*}+\frac{\partial u}{\partial h_{t-1}^{*}}$
$=\Lambda\left(z_{t}^{*},x_{t}\right)+\frac{\partial\Lambda}{\partial z_{t}^{*}}\frac{\partial\varphi}{\partial h_{t-1}^{*}}\mathrm{diag}\left(h_{t-1}^{*}\right)+\frac{\partial u}{\partial z_{t}^{*}}\frac{\partial\varphi}{\partial h_{t-1}^{*}}\,,$
where we used the implicit dependency of $z_{t}^{*}$ on $h_{t-1}^{*}$ via $\varphi$ to differentiate $\Lambda$ and $u$. ∎
[…]
Figure 7: State-to-state Jacobian according to Equation 10 of a single layer implicit Mamba2. To simplify the visualization, only one state variable of each head is plotted. Left: Jacobian computed with automatic differentiation through the fixed point iteration. Center: Jacobian computed with Equation 10. Right: Relative difference $\frac{\lvert J_{\text{autograd}}-J_{\text{theorem}}\rvert}{\lvert J_{\text{autograd}}\rvert+\lvert J_{\text{theorem}}\rvert+\epsilon}$. Top: Randomly initialized model. Bottom: Trained model.
Appendix B Algebraic Structure of Finite State Machines
@@
This section provides a basic introduction to the word problem and its relation to simulating finite state machines (FSMs). We start with some results in the circuit complexity and then relate them to the properties of FSMs.
B.1 Circuit Complexity
@@
B.2 Algebraic Concepts of Finite State Machines
@@
Monoid where each element has a a corresponding inverse element are called groups.
@@
Appendix C Relaxed Non-Solvable Word Problems
[…]
  1. 1.
    Which density of entities is required to learn the intrinsic algorithm of the state tracking task?
  2. 2.
    How many self-iterations do our implicit models need to conduct to learn the intrinsic algorithm?
1 Yesterday Mary travelled to the kitchen.
2 Fred went to the park yesterday.
3 This morning Fred journeyed to the bedroom.
4 Bill went back to the bedroom yesterday.
5 Q: Where was Fred before the bedroom?
1 Yesterday Bill went to the school.
2 Mary went to the cinema yesterday.
3 This morning Julie went back to the kitchen.
4 Yesterday Julie went to the office.
5 Q: Where was Julie before the kitchen?
Table 2: Two examples of tracking entities in natural language from the catbAbi dataset (Schlag et al., 2021).
Figure 8: Examples of word problems with mixed solvable and non-solvable tokens. We refer to tokens from a non-solvable group as non-solvable tokens, and to tokens from a star free language as solvable tokens.
@@ D.2 CatbAbI
The models, both implicit and explicit, comprise up to three layers of the Mamba2 architecture with an embedding dimension of 256, a state dimension of 16 (expansion factor 2), and a head dimension of 32. We trained the models using batch sizes of 128 and 256, and learning rates of 0.0001, 0.0005, 0.001, and 0.005. The models were trained for 15,000 steps, with the implicit model specifically trained for 5,000 steps in unrolling mode, utilizing 32 steps with normal gradient checkpointing, followed by 10,000 steps of self-iteration fixed-point search. The self-iteration included a stop threshold of 0.03 and a training and testing maximum number of steps 50 and 200 , respectively, and phantom gradient parameters of 6 steps with ($\lambda=0.5$). Data were packed in sequences of length 200 tokens as per (Schlag et al., 2021). Figures 10 and 11 illustrate the validation accuracy of the explicit and implicit Mamba2 models on the CatbAbI dataset, respectively. Additionally, Figure 12 plots the number of iterations required for the implicit model to reach a fixed point on the validation set of the CatbAbI dataset.
@@ D.3 Language Modeling
Appendix E Additional Results
E.1 Inference Dynamics and Convergence of Implicit Language Models
Model D-Pile Perplexity Inference Steps Rel. Diff.
130M Mamba2(4+1) 13.76 14 0.035
Mamba2(24+4) 12.86 62 0.036
Llama(4+1) 12.73 13 0.014
Llama(32+4) 11.73 53 0.033
350M Mamba2(4+1) 10.02 10 0.012
Mamba2(24+4) 9.70 49 0.024
Llama(4+1) 9.66 12 0.015
Llama(32+4) 9.43 57 0.037
760M Mamba2(4+1) 8.60 10 0.013
Mamba2(24+4) 8.35 45 0.025
Llama(4+1) 8.27 12 0.014
Llama(32+4) 7.90 77 0.044
1.3B Mamba2(4+1) 7.97 10 0.013
Mamba2(24+4) 7.70 47 0.029
Llama(4+1) 7.66 13 0.015
Llama(32+4) 7.24 69 0.048
E.2 Length Extrapolation Capabilities in Pretrained Models
Model 2048 4096 8192 16384
$\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$
130M Mamba2 13.72 13.44 13.92 14.91
Mamba2 13.05 12.72 12.76 12.94
Mamba2(4+1)-ours 13.76 13.04 13.06 13.25
Mamba2(24+4)-ours 12.86 12.53 12.55 12.78
370M Mamba2 10.55 10.28 11.20 13.89
Mamba2 10.18 9.94 10.07 10.34
Mamba2(4+1)-ours 10.02 9.74 9.73 9.86
Mamba2(24+4)-ours 9.70 9.45 9.44 9.59
760M Mamba2 9.23 9.45 34.99 231.25
Mamba2 8.98 8.94 10.38 12.56
Mamba2(4+1)-ours 8.60 8.38 8.40 8.55
Mamba2(24+4)-ours 8.35 8.16 8.18 8.33
1.3B Mamba2 8.40 8.47 12.69 25.37
Mamba2 8.28 8.24 9.96 15.06
Mamba2(4+1)-ours 7.97 7.81 8.36 9.53
Mamba2(24+4)-ours 7.70 7.59 8.19 9.63
E.3 Effective Duality between Simultaneous Mode and Sequential Mode in SSMs
Ground Truth Simultaneous Sequential